-2x-9y=-25
-4x-9y=-23
How do you solve this problem

Answer by mananth(16949)

-2x-9y= -25 .............1
-4x-9y =-23 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
2x+9y= 25
-4x-9y= -23
Add the two equations
-2x= 2
/ -2
x = -1
plug value of x in (1)
-2 x + -9 y = -25
2 + -9 y = -25
-9 y = -25-2
-9 y = -27
y = 3

m.ananth@hotmail.ca